OpenStudy (anonymous):

f(x)=x^3-6x; g(x)=3x...find (g o f)(x)

OpenStudy (anonymous):

ok gf(x)=3x^3-18x

OpenStudy (anonymous):

(g o f)(x)=g(f(x))=g(x^3-6x)=3(x^3-6x)=3x^3-18x

OpenStudy (anonymous):

gof(x)=3x^3-18x

OpenStudy (anonymous):

thanks guys